### v1

@@ Abstract
Since the launch of Uniswap and other AMM protocols, the DeFi industry has evolved from simple constant product functions with uniform liquidity distribution across the entire price axis to more advanced mechanisms that allow Liquidity Providers (LPs) to concentrate capital within selected price ranges. This evolution has introduced new research challenges focused on optimizing capital allocation in Decentralized Exchanges (DEXs) under dynamic market conditions. In this paper, we present a methodology for finding optimal liquidity provision strategies in DEXs within a specific family of $\tau$-reset strategies. The approach is detailed step by step and includes an original method for approximating historical liquidity within active pool ranges using a parametric model that does not rely on historical liquidity data. We find optimal LP strategies using a machine learning approach, evaluate performance over an out-of-time period, and compare the resulting strategies against a uniform benchmark. All experiments were conducted using a custom backtesting framework specifically developed for Concentrated Liquidity Market Makers (CLMMs). The effectiveness and flexibility of the proposed methodology are demonstrated across various Uniswap v3 trading pairs, and also benchmarked against an alternative backtesting and strategy development tool.
Keywords:
Decentralized exchange Automated market-making Liquidity provision Uniswap Decentralized finance Blockchain.
footnotetext: E-mail addresses: tapwi93@gmail.com (A. Urusov), rostislavberezovskiy@gmail.com (R. Berezovskiy), acryptokrestenko@gmail.com (A. Krestenko), kornilov.ag94@gmail.com (A. Kornilov).
@@ 1 Introduction
Decentralized Finance (DeFi) has become one of the fastest-growing areas of the blockchain industry, offering an open, programmable architecture to create, execute, and automate financial services through smart contracts, which are self-executing programs that run on-chain without intermediaries [1]. A central component of this ecosystem is the Decentralized Exchange (DEX), which is based on the Automated Market Maker (AMM) mechanism. In an AMM, liquidity is pooled and trading prices are set algorithmically rather than through a traditional order book. Since the launch of Uniswap V1/V2 [2], such protocols have provided a simple and accessible trading model in which pricing is governed by a predefined mathematical rule, most notably the constant product function [3].
@@
Consequently, optimal liquidity provision has become a critical research objective. Active liquidity management requires LPs to make decisions that weigh current and expected market conditions, risk exposure, and competition within pools. Therefore, designing effective liquidity management strategies requires a deep analysis of pool dynamics and the development of specialized tools to test research hypotheses.
This paper introduces a general methodology for developing models that support optimal liquidity provision decisions on CLMM platforms. At the core of the approach is a custom-designed backtesting framework that approximates historical liquidity from pool swap data, eliminating the need for historical in-range liquidity data. The resulting tool is lightweight, scalable, and suitable not only for LP strategies but also for analyzing prospective AMM designs. The methodology comprises:
  • Approximating historical liquidity within the active ranges of the pool.
  • Formalizing the LP reward optimization problem.
  • Training machine learning models with market features.
  • Evaluating the models on an out-of-time period and comparing to a uniform benchmark strategy.
  • Benchmark comparison against an alternative DeFi backtesting and strategy development tool.
@@
The remainder of the paper is organized as follows. Section 2 formalizes the problem and describes key concepts, strategy structure, overall ML architecture, and reward modeling techniques. Section 3 reviews related work. Section 4 presents an empirical study of historical Uniswap V3 pool data, covering liquidity reconstruction, model training, and performance evaluation. Section 5 summarizes the findings and provides directions for future research.
2 Background and Proposed Solution
This section presents a methodology for determining the optimal liquidity provision strategy using historical data analysis. The overall approach is based on modeling the processes inside a liquidity pool and involves the following consecutive stages: (i) modeling pool dynamics over a historical period, (ii) identifying the optimal LP strategies, and (iii) developing and testing an ML model for future decision making. Practical results, modeling details of the ML implementation, and a comparison with an alternative toolkit are given in Section 4.
To answer the question of what the optimal strategy should be in real time under current market conditions, the abstract ML model must be trained on historical data. The market state at each capital allocation moment is described by a set of characteristics (features), and the target variables are vectors of optimal capital shares in the pool price ranges. The vector of optimal capital shares constitutes the historical optimal strategy. Determining such an optimal strategy requires a strictly formalized methodology that integrates conceptual, architectural, and mathematical tools. We present the complete methodology, including the main concepts and research tools utilized in this work.
@@ 2.1 Key Concepts and External Data
After choosing the trading pair (e.g. USDC/ETHIn this research, ETH denotes Wrapped ETH (WETH) — the ERC-20 tokenized version of Ether that is used for interacting with AMM protocols.) and the specific DEX pool with a given fee tier $\Gamma$ for which the optimal liquidity model will be constructed, we select a historical time interval—the modeling period—over which the simulation is carried out. Next, we extract the set of swap transactions executed in the selected pool during this period and reconstruct the in-pool prices from those swaps. The resulting set of historical prices is denoted $\bm{P}=\{p_{0},\,p_{1},\,\dots,\,p_{H}\}$. Price intervals that contain elements of $\bm{P}$ are referred to as active (or working) ranges.
@@
2.2 Baseline Strategy Type
@@
At the moment that the LP allocates capital $W^{\mathrm{LP}}$ to the pool as liquidity $L^{\mathrm{LP}}$, as in [5] we identify reference bucket $B_{M}\in\beta$ with the index $M\in[1,N]$ that contains the initial price $p_{0}\in\bm{P}$. The parameter $\tau$ is then fixed: It specifies how many buckets on either side of the reference bucket will also receive LP liquidity. These buckets (and their associated price ranges) are called the LP-liquid buckets and are denoted by set $\beta^{\mathrm{LP}}_{L}=\{B_{M-\tau},\ldots,B_{M+\tau}\}$ with size $12\tau+$ 1. The core rebalancing rule is as follows: If the next observed price falls into a bucket with index $M\pm\vartheta$ where $\vartheta>\tau$, the reference index $M$ is updated and the liquidity of the LP is reallocated to the new set of buckets, as illustrated in Fig. 1.
Figure 1: Visualization of the LP-liquidity reallocation process under the $\tau$-reset strategy ($\tau=1$) on a fixed bucket partition $\beta$ over seven epochs.
@@ 2.3 Concept of Liquidity
    Depending on the current pool price $p$ relative to the placement interval $[p_{a_{i}},\,p_{b_{i}}]$, the capital $W^{\mathrm{LP}}_{i}$ is converted into the real reserves $x_{i}$ and $y_{i}$ of bucket $i$, corresponding to the two pool tokens $A$ (e.g. ETH) and $B$ (e.g. USDC), such that
@@
    These reserves reside inside bucket $B_{i}$ as liquidity $L_{i}$. The fundamental invariant Uniswap v3 that links $L_{i}$ to the reserves $(x_{i},y_{i})$ within $[p_{a_{i}},p_{b_{i}}]$ is
@@
    In this case, the allocation of capital $W_{i}^{\mathrm{LP}}$ proceeds by computing the reserves real of tokens $A$ and $B$ that correspond to the bucket $B_{i}$, given its price interval $[p_{a_{i}},p_{b_{i}}]$ and the price in the range $p$. In contrast, in the inverse problem, when the desired real reserves $x_{i}>0$ and $y_{i}>0$ are specified in advance but the partition $\beta$ is not, the limits of the bucket $[p_{a_{i}},p_{b_{i}}]$ can be recovered analogously.
@@
[…]
@@ 2.4 Problem Statement
[…]
@@
[…]
@@
The shares $\bm{r}_{e_{i}}^{\mathrm{LP}}$ are derived from the liquidity vector $\bm{L}_{e_{i}}^{\mathrm{LP}}$ and an epoch-specific weight vector $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}=\bigl{(}\alpha_{1}^{\mathrm{LP}},\alpha_{2}^{\mathrm{LP}},\ldots,\alpha_{N}^{\mathrm{LP}}\bigr{)}$, $\sum_{n=1}^{N}\alpha_{n}^{\mathrm{LP}}=1$, $\alpha_{n}^{\mathrm{LP}}\geq 0$, where $w_{n}^{\mathrm{LP}}=\alpha_{n}^{\mathrm{LP}}\,W_{e_{i}}^{\mathrm{LP}}$. The $\tau$-reset strategy $\varphi^{\tau}$ imposes sparsity on $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}$: positive weights appear only in the $(2\tau+1)$ buckets centered on the reference index $M$. Hence
@@
2.5 Architecture of the Core ML Model
@@
[…]
@@
where $w_{1},w_{2},w_{3}>0$ and $w_{1}+w_{2}+w_{3}=1$. The weights are chosen to minimize the error on the train date. A schematic of the ensemble is shown in Fig. 2.

### v2

@@ Abstract
Concentrated Liquidity Market Makers (CLMMs) represent a fundamental innovation in market microstructure, transforming liquidity provision from passive portfolio allocation to active risk management. This evolution creates significant challenges for performance evaluation and strategy optimization, particularly due to the absence of comprehensive historical liquidity data. We address these challenges through a novel methodological framework that reconstructs historical liquidity states from swap transaction data, enabling rigorous backtesting of dynamic liquidity provision strategies. Our parametric reconstruction method achieves high accuracy (approximation errors averaging around 2%) without relying on historical liquidity snapshots, addressing a critical data gap in decentralized finance research. We apply this framework to evaluate tau-reset strategies–dynamic liquidity reallocation approaches that respond to market movements–across multiple Uniswap v3 pools. Using machine learning to optimize strategy parameters based on market conditions, we identify consistent outperformance (13–23% higher fees) compared to uniform allocation benchmarks. Our analysis reveals important insights into the risk-return tradeoffs in automated market making, including the critical role of impermanent loss as a dominant risk factor and the effectiveness of asymmetric strategy modifications for capital preservation. These findings contribute to the broader understanding of market microstructure in decentralized exchanges, providing both methodological innovations for performance evaluation and practical insights for liquidity providers navigating this evolving financial landscape.
footnotetext: E‑mail addresses: tapwi93@gmail.com (A. Urusov), rostislavberezovskiy@gmail.com (R. Berezovskiy), acryptokrestenko@gmail.com (A. Krestenko), kornilov.ag94@gmail.com (A. Kornilov), y.yanovich@skoltech.ru (Y. Yanovich).
@@ 1 Introduction
Decentralized Finance (DeFi) has become one of the fastest-growing areas of the blockchain industry, offering an open, programmable architecture to create, execute, and automate financial services through smart contracts, which are self‑executing programs that run on-chain without intermediaries [1]. The DeFi ecosystem has evolved rapidly to include decentralized exchanges, lending protocols, and derivative markets [33]. A central component of this ecosystem is the Decentralized Exchange (DEX), which is based on the Automated Market Maker (AMM) mechanism. In an AMM, liquidity is pooled and trading prices are set algorithmically rather than through a traditional order book. Since the launch of Uniswap V1/V2 [2], such protocols have provided a simple and accessible trading model in which pricing is governed by a predefined mathematical rule, most notably the constant product function [3]. This represents a departure from traditional order-book exchanges and has stimulated significant research into decentralized exchange mechanisms [34].
@@
Consequently, optimal liquidity provision has become a critical research objective. However, developing and evaluating such strategies presents significant challenges. A key difficulty is the absence of comprehensive historical liquidity data needed for accurate backtesting. While swap transaction histories are typically available, the precise distribution of liquidity across price ranges at historical moments is often inaccessible or incomplete. This data gap hinders the development of data-driven approaches to liquidity optimization. Active liquidity management requires LPs to make decisions that weigh current and expected market conditions, risk exposure, and competition within pools. Therefore, designing effective liquidity management strategies requires a deep analysis of pool dynamics and the development of specialized tools to test research hypotheses.
To address these challenges, this paper introduces a general methodology for developing models that support optimal liquidity provision decisions on Concentrated Liquidity Market Makers (CLMM) platforms. The core innovation lies in reconstructing historical liquidity states from readily available swap data, enabling rigorous strategy development without dependency on historical liquidity snapshots. This paper makes three primary contributions to the financial analysis literature:
  1. 1.
    Methodological Innovation: We develop and validate a parametric methodology for reconstructing historical liquidity profiles solely from swap transaction data, eliminating dependency on historical liquidity snapshots. This approach enables accurate performance evaluation and strategy backtesting even when complete historical data is unavailable.
  2. 2.
    Empirical Strategy Analysis: We implement a comprehensive framework for developing and evaluating dynamic $\tau$-reset strategies across multiple asset pairs and market conditions. Our analysis identifies consistent patterns in optimal strategy parameters and performance characteristics, providing actionable insights for liquidity providers.
  3. 3.
    Financial Market Insights: Beyond methodological contributions, our findings offer important insights into market microstructure in decentralized exchanges. We demonstrate how liquidity provision strategies affect risk-return tradeoffs, how market conditions influence optimal strategy parameters, and how modified strategy designs can mitigate specific risks inherent in CLMM participation.
@@
The remainder of this paper is organized as follows. Section 2 provides the necessary background on CLMM mechanics, $\tau$-reset strategies, Section 3 describes the related work. Section 4 formalizes the problem and describes our proposed methodology, including liquidity approximation, reward modeling, and machine learning architecture. Section 5 presents an empirical study of historical Uniswap V3 pool data, covering liquidity reconstruction, model training, performance evaluation, and comparison with alternative tools. Section 6 discusses theoretical and practical implications, and Section 7 summarizes the findings and provides directions for future research.
2 Background
This section provides the necessary background for understanding our proposed methodology. We first review key concepts and terminology related to concentrated liquidity markets, then discuss the $\tau$-reset strategy family that forms the basis of our approach, and finally survey related work in the field.
@@ 2.1 Key Concepts and External Data
After choosing the trading pair (e.g. USDC/ETHIn this research, ETH denotes Wrapped ETH (WETH) — the ERC‑20 tokenized version of Ether that is used for interacting with AMM protocols.) and the specific DEX pool with a given fee tier $\Gamma$ for which the optimal liquidity model will be constructed, we select a historical time interval—the modeling period—over which the simulation is carried out. Next, we extract the set of swap transactions executed in the selected pool during this period and reconstruct the in-pool prices from those swaps. The resulting set of historical prices is denoted $\bm{P}=\{p_{0},\,p_{1},\,\dots,\,p_{H}\}$. Price intervals that contain elements of $\bm{P}$ are referred to as active (or working) ranges.
@@
2.2 Baseline Strategy Type: $\tau$-Reset Strategies
@@ 2.2.1 Bucket partition
At the moment that the LP allocates capital $W^{\mathrm{LP}}$ to the pool as liquidity $L^{\mathrm{LP}}$, as in [5] we identify reference bucket $B_{M}\in\beta$ with the index $M\in[1,N]$ that contains the initial price $p_{0}\in\bm{P}$. The parameter $\tau$ is then fixed: It specifies how many buckets on either side of the reference bucket will also receive LP liquidity. These buckets (and their associated price ranges) are called the LP-liquid buckets and are denoted by set $\beta^{\mathrm{LP}}_{L}=\{B_{M-\tau},\ldots,B_{M+\tau}\}$ with size $2\tau+1$. The core rebalancing rule is as follows: If the next observed price falls into a bucket with index $M\pm\vartheta$ where $\vartheta>\tau$, the reference index $M$ is updated and the liquidity of the LP is reallocated to the new set of buckets, as illustrated in Fig. 1.
Figure 1: Visualization of the LP-liquidity reallocation process under the $\tau$-reset strategy ($\tau=1$) on a fixed bucket partition $\beta$ over eight epochs.
@@ 2.3 Concept of Liquidity
    Depending on the current pool price $p$ relative to the placement interval $p_{a_{i}},\,p_{b_{i}}]$, the capital $W^{\mathrm{LP}}_{i}$ is converted into the real reserves $x_{i}$ and $y_{i}$ of bucket $i$, corresponding to the two pool tokens $A$ (e.g. ETH) and $B$ (e.g. USDC), such that
@@
    These reserves reside inside bucket $B_{i}$ as liquidity $L_{i}$. The fundamental invariant Uniswap v3 that links $L_{i}$ to the reserves $(x_{i},y_{i})$ within $p_{a_{i}},p_{b_{i}}]$ is
@@
    In this case, the allocation of capital $W_{i}^{\mathrm{LP}}$ proceeds by computing the reserves real of tokens $A$ and $B$ that correspond to the bucket $B_{i}$, given its price interval $p_{a_{i}},p_{b_{i}}]$ and the price in the range $p$. In contrast, in the inverse problem, when the desired real reserves $x_{i}>0$ and $y_{i}>0$ are specified in advance but the partition $\beta$ is not, the limits of the bucket $p_{a_{i}},p_{b_{i}}]$ can be recovered analogously.
@@
    $\mathcal{V}_{i}\!\left(L_{i},p,B_{i}\right)=\begin{cases}\bigl(L_{i}\,\Delta^{x_{i}}_{p_{a_{i}},p_{b_{i}}},\,0\bigr),&p\leq p_{a_{i}},\\[4.0pt] \bigl(L_{i}\,\Delta^{x_{i}}_{p,p_{b_{i}}},\,L_{i}\,\Delta^{y_{i}}_{p_{a_{i}},p}\bigr),&p_{a_{i}}<p<p_{b_{i}},\\[4.0pt] \bigl(0,\,L_{i}\,\Delta^{y_{i}}_{p_{a_{i}},p_{b_{i}}}\bigr),&p\geq p_{b_{i}}.\end{cases}$ (6)
@@
3 Related Work
This subsection reviews existing literature on liquidity provision in CLMM markets, covering specific strategies, backtesting frameworks, and machine learning applications. Recent systematic reviews compare the architectures and performance of centralized versus decentralized exchanges (CEX and DEX) [44]. Beyond core exchange mechanics, DeFi protocols are enabling novel financial interactions in adjacent sectors such as GameFi, where liquidity provision and staking mechanisms are integral to ’play-to-earn’ economic models [45].
The concept of liquidity provision in concentrated liquidity markets has been studied from multiple perspectives, including impermanent loss, trading fee dynamics, DEX market design, and liquidity provision strategies.
The option-like properties of liquidity provision in both CPMM and CLMM settings have been thoroughly analyzed, with foundational work on token reserves, swaps, fees, and liquidity calculations provided in [7, 9]. Building on this, [10] developed a methodology to estimate accumulated fees in liquidity pools under basic price behavior assumptions, though without explicit consideration of trading volumes.
From a trader’s perspective, [11] compares price discovery processes in DEXs versus centralized markets and studies the behavior of informed traders. Meanwhile, [12] analyzes bid-ask prices in liquidity pools and derives the value of no-trade gaps. Arbitrage opportunities in DEXs have been systematically studied, including simple arbitrage strategies [40] and more complex cyclic arbitrage patterns [41].
Research on impermanent loss includes [13], which studies LP risks and potential returns, providing numerical estimations from real-world data. [14] offers the concerning estimate that approximately half of LPs lose money in their dataset, highlighting the complexity of concentrated liquidity provision. Further analysis by [35] examines the risk profile of liquidity providers, while [36] and [37] provide additional theoretical frameworks for understanding impermanent loss in both traditional and concentrated liquidity settings. [15] provides concise results for estimating impermanent loss and generalizes the concept to allow comparison with various benchmark strategies.
Hedging impermanent loss with traditional options has been explored in [16], which applies arbitrage-based methods to produce both static and dynamic hedges under Black-Scholes-Merton and log-normal stochastic volatility models.
The $\tau$-reset strategy family, central to our work, builds on concepts introduced in [5, 6]. Related strategy research includes [17], which considers accumulated fees, rebalancing costs, and liquidity concentration risk in continuous-time models for specific market conditions. [18] models liquidity dynamics in CLMM pools in continuous time with rebalancing requirements, while [19] approaches liquidity provision as an optimal stopping problem.
Machine learning applications in CLMM markets are gaining traction. [20] applies deep reinforcement learning to maximize accumulated fees while considering loss-versus-rebalancing risk, optimizing price interval parameters. Similarly, [21] demonstrates superior strategy performance compared to benchmarks using deep reinforcement learning on ETH/USDC and ETH/USDT pools. Beyond strategy optimization, [22] applies reinforcement learning to AMM design itself, aiming to reduce impermanent loss and slippage.
Market design considerations are addressed in [23], which introduces generalized decentralized liquidity pools with dynamic bonding curves, and [24], which extends the problem by considering the trading platform as a player competing for fees with LPs.
Competition among LPs has been modeled using mean-field games in [25], analyzing resulting liquidity distributions and incorporating just-in-time liquidity. Beyond market microstructure risks, broader DeFi protocol risks including smart contract vulnerabilities, regulatory uncertainty, and systemic risks have been examined in the literature [42, 43]. A specific focus on just-in-time liquidity appears in [26].
Finally, backtesting frameworks for CLMM strategies are essential for practical validation. Our previous work [6] provides the foundation for the backtesting approach extended in this paper. The current research applies accumulated knowledge of concentrated liquidity mechanics to extend this framework and enable data-driven liquidity provision optimization using machine learning methods within the $\tau$-reset strategy family.
4 Problem Statement and Proposed Solution
This section formalizes the liquidity provision optimization problem and presents our proposed methodology for developing optimal $\tau$-reset strategies using historical data analysis.
@@ 4.1 Problem Statement
where $\cdot$$_{+}$ is applied to the change of reserves $\Delta\mathcal{V}^{q}_{n}$ and $\Gamma$ denotes the pool’s fee tier. After this, let $\bm{m}_{e_{i}}=(\,p_{\gamma_{i}},\,1\,)$ be the vector of token prices expressed in the chosen numéraire at the epoch-closing price $p_{\gamma_{i}}$. Assuming efficient on-chain trading so that in-pool prices are close to market prices, the market worth of the fees earned by the pool during epoch $e_{i}$ is
@@
with $\bm{m}_{e_{i}}=(p_{\gamma_{i}},\,1)$. Aggregating over all epochs yields
@@
The shares $\bm{r}_{e_{i}}^{\mathrm{LP}}$ are derived from the liquidity vector $\bm{L}_{e_{i}}^{\mathrm{LP}}$ and an epoch-specific weight vector $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}=\bigl(\alpha_{1}^{\mathrm{LP}},\alpha_{2}^{\mathrm{LP}},\ldots,\alpha_{N}^{\mathrm{LP}}\bigr)$, $\sum_{n=1}^{N}\alpha_{n}^{\mathrm{LP}}=1$, $\alpha_{n}^{\mathrm{LP}}\geq 0$, where $w_{n}^{\mathrm{LP}}=\alpha_{n}^{\mathrm{LP}}\,W_{e_{i}}^{\mathrm{LP}}$. The $\tau$-reset strategy $\varphi^{\tau}$ imposes sparsity on $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}$: positive weights appear only in the $(2\tau+1)$ buckets centered on the reference index $M$. For the uniform benchmark strategy, all non-zero elements of $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}$ are equal to $\frac{1}{2\tau+1}$. Hence
@@
4.2 Machine Learning Architecture
@@
For each epoch $e_{i}$, the optimal LP strategy is represented by the weight vector $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}\bigl(\varphi_{e_{i}}^{\mathrm{OPT}}\bigr)\in\mathbb{R}^{N}_{+}$, whose non-zero entries $2\tau+1$ occupy the buckets $j\in[M-\tau,\,M+\tau]$. Collect these entries in the vector $\bm{\rho}_{e_{i}}^{\mathrm{LP}}=(\alpha_{M-\tau}^{\mathrm{LP}},\ldots,\alpha_{M}^{\mathrm{LP}},\ldots,\alpha_{M+\tau}^{\mathrm{LP}})\in\mathbb{R}^{2\tau+1}_{+}$, and exploit the $\tau$-symmetry $\alpha_{M-\tau+k}^{\mathrm{LP}}=\alpha_{M+\tau-k}^{\mathrm{LP}}$ for $k\in[0,\,\tau-1]$ to define the reduced representation $\hat{\bm{\rho}}_{e_{i}}^{\mathrm{LP}}=(\alpha_{M}^{\mathrm{LP}},\,2\alpha_{M+1}^{\mathrm{LP}},\,\ldots,\,2\alpha_{M+\tau}^{\mathrm{LP}})\in\mathbb{R}_{+}^{\tau+1}$. Stacking the epoch vectors yields the target vector
@@
where $w_{1},w_{2},w_{3}>0$ and $w_{1}+w_{2}+w_{3}=1$. The weights are chosen to minimize the error on the train date. The ensemble leverages the architectural heterogeneity of MLPs, CatBoost and LSTMs to capture a broad spectrum of signal types—from static tabular relationships to non-linear interactions and latent compositional structures—within a unified learning framework. A schematic of the ensemble is shown in Fig. 2.